OpenStudy (anonymous):

how do you do 3x-4 = 3(2x-1)+7

OpenStudy (bahrom7893):

3x-4 = 6x - 3 +7 3x - 6x = 8 -3x = 8 x = -8/3

OpenStudy (anonymous):

thanks